OpenStudy (anonymous):

Consider the function y= -2-cos(x-pi). What effect does pi have on the basic graph? a. Vertical shift down pi units b. Horizontal shift left pi units c. Vertical stretch by factor pi d. Horizontal shift right pi units

OpenStudy (anonymous):

y= K +A cos ( Bx + h )

OpenStudy (anonymous):

so shift left pi units?

OpenStudy (anonymous):

nooooooo to the Right

OpenStudy (anonymous):

take x=0

OpenStudy (anonymous):

try it on y= -2-cos(x)

OpenStudy (anonymous):

then try it on y= -2-cos(x-pi)

OpenStudy (anonymous):

you will get for the 1st one : -3 for the 2nd one : -1

OpenStudy (anonymous):

so its moving to the RIGHT

OpenStudy (anonymous):

@kaykim95 did you get it??